Find five solution for x and y for 2(x+1)=3(y-1)
Answer by vidya pattar(14)

Solution:
2(x+1)= 3( y - 1)
use distributive property"
2x + 2 = 3y - 3
add 3 on both sides we get:
2x + 2 + 3= 3y - 3 + 3
3y = 2x + 5
divide both sides by 3 we get '
y = ( 2x + 5 ) / 3
now take x = 0
then y = 5/3 , so solution 1) ( 0 , 5/3)
take x = 2
then y = ( 2*2 + 5 ) / 3 = 9/3 = 3
so the solution 2) = ( 2,3)
take x= -1
then y = ( 2 * -1 + 5 ) / 3 = 3/3= 1
hence the solution 3) is ( -1 , 1 )